OpenStudy (marigirl):

y=-(x-2)^2 + 4 . If I make x the subject, how will my function look like?

OpenStudy (anonymous):

you should expand out the (x-2) first. so what you have is: y = -[(x-2)(x-2)] + 4 what do you get?

OpenStudy (marigirl):

is it wrong to go: \[y=-(x-2)^{2} +4\] move the 4 on to the other side, \[y-4 = -(x-2)^{2}\]

OpenStudy (marigirl):

Hi Nuk64, I expanded (x-2) and collected like terms and got to \[-x ^{2}+4x=y\] .. im a bit confused how to go from there to make x the subject

OpenStudy (anonymous):

No, you could do that. It's not wrong. Then what you'd do is remove the negative sign. so you have: -(y-4) = (x-2)^2 then get rid of the square by taking the square root of each size.

OpenStudy (anonymous):

actually, your way's easier.... don't bother expanding.

OpenStudy (anonymous):

yep! and then all you have to do is add 2, and you get your function! :)

OpenStudy (marigirl):

oh yea cool, then i got \[x= \sqrt{-y+4} +2\]

OpenStudy (marigirl):

thank you! :D

OpenStudy (anonymous):

no worries. :)